1/2

Heinrich Heine,  Heinrich Heine,  Heinrich Heine,
The Lorelei  Lorelej’  The Lorelei
tradukita de L. W. Garnham  tradukita de Leopold Elb tradukita de Mark Twain

Tiu traduko estas poste iom modifita de aliuloj, ne estas la origina traduk-versio de Leopold Elb.

An ancient legend of the Rhine

I do not know what it signifies. That I am so sorrowful? A fable of old Times so terrifies, Leaves my heart so thoughtful.

Ne scias mi, kial mi estas Malgaja en la kor’. Antikva fabelo restas Por ĉiam en mia memor’.

I cannot divine what it meaneth, This haunting nameless pain: A tale of the bygone ages Keeps brooding through my

brain:

The air is cool and it darkens, And calmly flows the Rhine; The summit of the mountain he-

arkens In evening sunshine line.

Vespero jam ekmalheliĝas, Fluadas la Rejn’ en trank-

vil’, La supro de l’ monto lu-

miĝas En vespersuna bril’.

The faint air cools in the glo-

aming, And peaceful flows the Rhine, The thirsty summits are drin-

king The sunset’s flooding wine;

The most beautiful Maiden ent-

rances Above wonderfully there, Her beautiful golden attire glan-

ces, She combs her golden hair.

Belega knabino jen tronas Kun ora juvelar’, La belan vizaĝon kronas Mirinde la ora harar’.

The loveliest maiden is sitting High-throned in yon blue air, Her golden jewels are shining, She combs her golden hair;

With golden comb so lustrous, And thereby a song sings, It has a tone so wondrous, That powerful melody rings.

Ŝi uzas kombilon el oro Kaj dume kantas ŝi. Mirige tuŝas al koro Sorĉiga la melodi’.

She combs with comb that is gol-

den, And sings a weird refrain That steeps in a deadly enchant-

ment The listener’s ravished brain:

The shipper in the little ship It effects with woe sad might; He does not see the rocky slip, He only regards dreaded height.

Kaj sopiregon eksentas Ŝipisto dum sia vetur’; Rifaron li ne priatentas, Al monto rigardas li nur.

The doomed in his drifting shal-

lop, Is tranced with the sad sweet to-

ne, He sees not the yawing breakers, He sees but the maid alone:

…  …  …

2/2

I believe the turbulent waves Swallow the last shipper and

boat; She with her singing craves All to visit her magic moat.

Mi kredas, ke dronos finfi-

ne Ŝipisto kun sia boat’, Kaj tion kaŭzis feine La Loreleja kantad’.

The pitiless billwos engulf him!- So perish sailor and bark; And this, with her baleful sin-

ging, Is the Lorelei’s gruesome work.

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Anglan de L. W. Garnham.

Arg-2-14 (2003-10-11 23:04:57)

L.W. Garnham, Bachelor of Arts, LEGENDS OF THE RHINE; mentioned in: A Tramp Abroad. Vol 1-2. Leibzig: Tauchnitz, 1880 Band I, Mark Twain 1880 Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en Esperanton de Leopold Elb (†1912-08.04).

Arg-2-71 (2005-02-03 19:06:30)

La traduko troviĝas sur paĝo 84 de la lernolibro �Post la kurso“ de Wilhelm kaj Hans Wingen, eldonita de Ludwig Pickel en Nurembergo (Nürnberg), Germanio.

Traduko de la Germana poemo “Die Lorelei” de Heinrich Heine (∗1797-12-13 – †1856-02-17) en la Anglan de Mark Twain.

Arg-2-11 (2003-10-13 04:42:59)

A Tramp Abroad. Vol 1-2. Leibzig: Tauchnitz, 1880 Band I, Mark Twain 1880 (rf. http: // www. loreley. com/ loreley/ marctwai. htm )

http://www.poezio.net/version?poem-id=2&version-id=14,71,11